OpenStudy (anonymous):

What is the volume of a cone that has a radius of 12 m and a height of 3 m? Use 3.14 to approximate pi and express your answer in hundredths. ___m3

OpenStudy (anonymous):

I got 66.978 m3

OpenStudy (cookielate):

V≈452.39m³

OpenStudy (cookielate):

V=πr2h 3=π·122·3 3≈452.38934m³

OpenStudy (cookielate):

solution ^

OpenStudy (anonymous):

@cookielate like u pic

OpenStudy (cookielate):

thanks

OpenStudy (anonymous):

V≈452.39m³ is the answer?

OpenStudy (cookielate):

yes i solved it already

OpenStudy (cookielate):

V=πr2h 3=π·122·3 3≈452.38934m³

OpenStudy (anonymous):

Ok thx c: can u help me with some more?

OpenStudy (cookielate):

sure how many more

OpenStudy (anonymous):

Bout 6 small ones...

OpenStudy (cookielate):

okay

OpenStudy (anonymous):

What is the exact volume of a sphere that has a radius of 7.5 ft? ___ pi ft3 I got 1767.15

OpenStudy (cookielate):

V≈1767.15ft³ your right :)

OpenStudy (cookielate):

5 more

OpenStudy (anonymous):

What is the volume of a sphere that has a diameter of 42 cm? Use 3.14 to approximate pi and express your answer in hundredths. ___cm3 I got 38792.39

OpenStudy (anonymous):

@cookielate

OpenStudy (cookielate):

V≈38792.39cm³

OpenStudy (cookielate):

your right

OpenStudy (anonymous):

Thx next one A cylindrical tank holds 14,130 in3 of water. Ida empties the tank into 25 identical smaller cylinders, each with a radius of 3 in. What is the height of the 25 smaller cylinders? Use 3.14 to approximate pi. ___in. I got 20

OpenStudy (cookielate):

V = pi * r^2 * h Plug in what we know: 14130 = 3.14 * 3^2 * h Simplify exponent: 14130 = 3.14 * 9 * h Multiply: 14130 = 28.26h Divide 28.26 to both sides: h = 500 Now divide by 25, since there are 25 cylinders. 500 / 25 = 20 So the height of each container is 20.

OpenStudy (cookielate):

your right

OpenStudy (cookielate):

your so good at this :)

OpenStudy (cookielate):

how many more do you need ?

OpenStudy (anonymous):

Thx but i just want to make sure my answers r right before I turn it in. Joseph is building a cone using modeling clay. The cone has a radius of 6 cm and a height of 12 cm. Joseph adds additional clay, keeping the radius the same, until the cone reaches a height of 18 cm. How much clay did Joseph add? Use 3.14 to approximate pi and express your final answer in hundredths. __cm3 I got 226.19

OpenStudy (anonymous):

I have 2 more after this one^

OpenStudy (cookielate):

Volume of a cone = ( pi x r squared x height ) divided by 3 (pi x 6 squared x 18) /3 - (pi x 6 squared x 12)/3 = extra clay 678.58 - 452.39 = 226.19 cm cubed

OpenStudy (anonymous):

A hard candy has a radius of 9 mm. The candy has a spherical soft center with a radius of 6 mm, with a hard shell surrounding it. What is the volume of the hard shell? Use 3.14 to approximate pi and express your final answer in hundredths. ___mm3 Im not sure on this one :/

OpenStudy (cookielate):

The answer is 141.3 because you first find the volume of the hard candy which is pi r squared then you subtract the middle of the candy

OpenStudy (anonymous):

Thx! c: last one!

OpenStudy (cookielate):

okay

OpenStudy (anonymous):

Which container holds the most? A. cylinder with radius 9 in. and height 11 in. B. sphere with radius 9 in. C. cone with radius 12 in. and height 20 in. I think A

OpenStudy (anonymous):

Oh wait now Im thinking B

OpenStudy (cookielate):

Cylinder: V = pi * r^2 * h V = 3.14 * 9^2 * 11 Simplify exponent: V = 3.14 * 81 * 11 Multiply: V = 2797.74 Sphere: V = 4/3 * pi * r^3 V = 4/3 * 3.14 * 9^3 Simplify exponent: V = 4/3 * 3.14 * 729 Multiply: V = 3052.08 Cone: V = 1/3 * pi * r^2 * h V = 1/3 * 3.14 * 12^2 * 20 Simplify exponent: V = 1/3 * 3.14 * 144 * 20 Multiply: V = 3014.4 So the Sphere holds the most.

OpenStudy (anonymous):

Thx!!!!! c:

OpenStudy (cookielate):

yup your right is b

OpenStudy (cookielate):

no problem anytime